# TIDMAD: Time Series Dataset for Discovering Dark Matter with AI Denoising

### v2

@@ 15. Guidelines:
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We publicly provide all of the data and code necessary to reproduce all experimental results of the paper. Please use the link in the abstract to access all necessary materials for reproduction.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
@@ 20. Guidelines:
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We publicly provide all of the data and code necessary to reproduce all experimental results of the paper. Please use the link in the abstract to access our data, benchmarking code, and analysis code.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
@@ 25. Guidelines:
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: In Section 2, we describe the data training splits. In section 3, we describe the hyperparameters used.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
@@ 30. Guidelines:
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: In Section 4, the second benchmark -- the dark matter limit -- is accompanied by the relevant confidence intervals and statistical significance of the scientific result. Additionally, the noise assumptions are discussed with respect to the denoising score.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
@@ 35. Guidelines:
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: As stated in the body of the paper, the model training and inference is conducted upon the SDSC Expanse cluster equipped with V100 GPUs. For benchmarking tasks, model training and inference over validation and science dataset takes less than O(100 GPU hours). The actual value varies slightly with model complexity and data size.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
@@ 40. Guidelines:
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We do not have human subjects or any data related to human beings. Safety: the authors do not foresee how this data could be directly or indirectly used to harm people. Security: the research does not introduce security vulnerabilities. Descrimination: our technology can not be used to discriminate people. Surveillance: we do not use surveillance data. Deception and Harassment: this is physics data -- it can not be used to facilitate deceptive interactions. Environment: our research does not promote negative impacts to the environment outside of the computational resources used for the research. Human Rights: no humans were used in this data collection or benchmark creation. Bias and fairness: our methods do not encode bias against certain people. Scientific bias is addressed in Section 5.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
@@ 45. Guidelines:
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: This is discussed in Section 5. This is no negative societal impact of the work performed.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
@@ 50. Guidelines:
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our data, models, and benchmarks relate to particle physics data. There is no risk for misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
@@ 55. Guidelines:
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The creators of the data are the authors of the paper. This data was produced by a particle physics detector operated by the authors.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The assets introduced in the paper are documented in the paper and provided with the publically available code repository.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: There was no crowdsourcing or human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: There were no human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: No LLM was used in the core methods of this research.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v3

@@ 1.2 Composition
    What data does each instance consist of? The data each instance consists of is a raw 8-bit integer from our digitizer. To convert the raw 8-bit integer to a voltage, each bit must be scaled by the ADC voltage i.e. multiply by $40mV/128$.
@@
    Is there a label or target associated with each instance? Yes, for the calibration data, each instance of the SQUID data corresponds to a target which is the instance in the SG data.
@@
    Is any information missing form individual instances? No.
@@
    Are relationships between individual instances made explicit? The instances are related because they come from the same detector just sampled at a different moment in time.
@@
    Are there recommended data splits? Yes, please see Section 2.l
@@
    Are there any errors, sources of noise, or redundancies in the dataset? There are no redundancies. Yes, there are many sources of detector noise.
@@
    Is the dataset self-contained, or does it link to or otherwise rely on external resources? The data is self-contained.
@@
    Does the dataset contain data that may be considered confidential? No
@@
    Does the dataset containd ata that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety? No.
Table 3: Summary of critical information about this data release. These are the data files used for training and benchmarking of the baseline algorithms provided.
Training Data Validation Data Science Data
File Name abra_training _00{00-19}.h5 abra_validation _00{00-19}.h5 abra_science _0{000-207}.h5
No. Data Points per File 2.01e9 2.01e9 4.01e9
HDF5 File Size 2.2 GB 2.2 GB 2.7 GB
ch1 Hardware Input SQUID SQUID SQUID
ch2 Hardware Input SG SG
Injected frequencies (Hz) [1100, 1200, … , 4.8M, 4.9M]
Injected amplitudes (mV) 50
Table 4: Summary of auxiliary files in this data release. These files provide an interesting challenge for the user, however were not used in the training or validation of the baseline models.
Aux Training Data Aux Validation Data
File Name abra_training_00{20-39}.h5 abra_validation_00{20-39}.h5
No. Data Points per File 2.01e9 2.01e9
HDF5 File Size 2.2 GB 2.2 GB
ch1 Hardware Input SQUID SQUID
ch2 Hardware Input SG SG
Injected Frequencies (Hz) [1100, 1200, … , 4.8M, 4.9M]
Injected Amplitudes (mV) 10
1.3 Collection process
  1. 1.
    How was the data associated with each instance acquired? The data associated with each instance is acquired by the ABRACADABRA detector. Full details can be viewed in Sections 1.1 and 2.
  2. 2.
    What mechanisms or procedures were used to collect the data? Full details can be viewed in Sections 1.1 and 2 and reference [28]. The hardware necessary for producing said data include, but are not limited to, an Oxford dilution refrigerator, 1T superconducting magnet, two-stage Magnicon SQUID, superconducting pickup loop, superconducting calibration loop, signal generator, digitizer, and data acquisition computer.
  3. 3.
    If the dataset is sampled from a larger set, what was the sampling strategy? The voltage produced by the SQUID and the SG are continuous. The instances are sampled from this continuous voltage stream at a constant rate of 10MS/s. The sampling strategy is deteministic and regular.
  4. 4.
    Who was involved in the data collection process and how were they compensated? To run the ABRACADABRA experiment, one graduate student, J. T. Fry, was needed. This graduate student was paid via NSF fellowship.
  5. 5.
    Over what timeframe was the data collected? The data were collected from 2/21/24 - 2/23/24.
  6. 6.
    Were any ethical review processes conducted? No. These data do not involve humans.
  7. 7.
    Does this dataset relate to people? No.
1.4 Preprocessing/cleaning/labeling
  1. 1.
    Was any preprocessing/cleaning/labeling of the data done? No.
1.5 Uses
  1. 1.
    Has the dataset been used for any tasks already? No, this dataset has not been used for any tasks yet.
  2. 2.
    Is there a repository that links to any or all papers or systems that use the dataset? No. This dataset has yet to be used outside of this paper.
  3. 3.
    What other tasks could the dataset be used for? As discussed in Section 5, these data can generally be used for training time series algorithms. Due to its high coherence and extensive length, it is perfect for cross cutting applications.
  4. 4.
    Is there anything about the composition of the dataset or the way it was collected and preprocessed that might impact future use? No.
  5. 5.
    Are there tasks for which the dataset should not be used? No.
1.6 Distribution
  1. 1.
    Will the dataset be distributed to third parties outside of the entity on behalf of which the dataset was created? Yes, the dataset is open to the public.
  2. 2.
    How will the dataset be distriuted? The dataset is publically available to be downloaded from the Open Science Data Federation cache. For download instructions, please see Section 2.
  3. 3.
    When will the dataset be distributed? The dataset is presently available.
  4. 4.
    Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)? No.
  5. 5.
    Have any third parties imposed IP-based or other restriction on the data associated with the instances? No.
  6. 6.
    Do any export controls or other regulatory restrictions apply to the dataset or to individual instances? No.
1.7 Maintenance
  1. 1.
    Who will be supporting/hosting/maintaining the dataset? The dataset is hosted at Open Science Data Federation (OSDF). The data storage at OSDF was offered to Dr. Aobo Li by the director of San Diego Supercomputer Center (SDSC). OSDF also provide distributed cache of the dataset across its global cache location. For more detail, please refer to OSDF Website***https://osg-htc.org/services/osdf.html. Dr. Aobo Li will be maintaining the dataset.
  2. 2.
    How can the owner/curator/mangager of the dataset be contacted? Please email Dr. Aobo Li at (liaobo77@ucsd.edu).
  3. 3.
    Is there an erratum? No.
  4. 4.
    Will the dataset be updated? No, the dataset will not be updated.
  5. 5.
    If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances? This dataset does not relate to people.
  6. 6.
    Will older versions of the dataset continue to be supported? Yes, they will continue to be supported.
  7. 7.
    If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so? No. Access to the ABRACADABRA detector is controlled.
Appendix 2 Dataset and code access
The data downloading and associated analysis scripts are available at https://github.com/jessicafry/TIDMAD. All data were uploaded to Open Science Data Federation and cached using their distributed cache system. The TIDMAD dataset can be downloaded using the download_data.py script provided in this GitHub repository. This script runs without any external dependencies. This script downloads data by generating a series of wget commands and executing them in a bash environment. download_data.py has the following argument:
  • --output_dir -o: Destination directory where the file will be downloaded, default: current working directory.
  • --cache -c: Which OSDF cache location should be used to download data. Options include [NY/NorCal/SoCal/Director(default)]:
    • NY: New York
    • NorCal: Sunnyvale
    • SoCal: San Diego
    • Director: automatically find the fastest cache location based on user’s location.
        WARNING: Director cache is sometimes unstable. We recommend switching to a different cache if the download fails.
  • --train_files -t: Number of training files to download, must be an integer between 0 and 20, default 20.
  • --validation_files -v: Number of validation files to download, must be an integer between 0 and 20, default 20.
  • --science_files -s: Number of science files to download, must be an integer between 0 and 208, default 208.
  • -f, --force: Directly proceed to download without showing the file size and asking the confirmation question.
  • -sk, --skip_downloaded: Skip the file that already exists at --output_dir.
  • -w, --weak: Download the weak signal version of training and validation files. In this version, the injected signal is 1/5 the amplitude of the normal version. This is a more challenging denoising task. Note that the normal version has a file range 0000-0019, while the weak version has a file range of 0020-0039.
  • -p, --print: Print out all wget commands instead of actually executing the download commands.
In the same github repository, we also provided a filelist.dat file which contains line-by-line wget command to download the entire dataset.An example wget command is given here:
We have also provided a link to download all trained models used to produce the results in the main manuscript. The models can be downloaded from https://drive.google.com/drive/folders/16ORX1b2zo1_lOYYAcRBgddBuYImj0Bxs?usp=share_link.
Appendix 3 Croissant metadata
We created a croissant metadata file TIDMAD_croissant.json using protocal presented in [4].
Appendix 4 Author statement
The authors of this paper all bear responsibility in the case of violation of rights. The information provided in the paper and supplementary material is truthful and accurate. The code from this paper is hosted, managed, and maintained by the paper author J. T. Fry at https://github.com/jessicafry/TIDMAD. The data from this paper is hosted, managed, and maintained by the paper author Dr. Aobo Li with download instructions in Section 2. The dataset is released under the Creative Commons Attribution (CC BY) license. The code is released under the GNU General Public License (GPL), version 3.